# MRT: Learning Compact Representations with Mixed RWKV-Transformer for Extreme Image Compression

### v1

@@ Abstract
Recent advances in extreme image compression have revealed that mapping pixel data into highly compact latent representations can significantly improve coding efficiency. However, most existing methods compress images into 2-D latent spaces via convolutional neural networks (CNNs) or Swin Transformers, which tend to retain substantial spatial redundancy, thereby limiting overall compression performance. In this paper, we propose a novel Mixed RWKV-Transformer (MRT) architecture that encodes images into more compact 1-D latent representations by synergistically integrating the complementary strengths of linear-attention-based RWKV and self-attention-based Transformer models. Specifically, MRT partitions each image into fixed-size windows, utilizing RWKV modules to capture global dependencies across windows and Transformer blocks to model local redundancies within each window. The hierarchical attention mechanism enables more efficient and compact representation learning in the 1-D domain. To further enhance compression efficiency, we introduce a dedicated RWKV Compression Model (RCM) tailored to the structure characteristics of the intermediate 1-D latent features in MRT. Extensive experiments on standard image compression benchmarks validate the effectiveness of our approach. The proposed MRT framework consistently achieves superior reconstruction quality at bitrates below 0.02 bits per pixel (bpp). Quantitative results based on the DISTS metric show that MRT significantly outperforms the state-of-the-art 2-D architecture GLC, achieving bitrate savings of $43.75\%$, $30.59\%$ on the Kodak and CLIC2020 test datasets, respectively.
Codehttps://github.com/luke1453lh/MRT
@@ Introduction
The explosive growth of visual data has intensified the demand for advanced image compression techniques capable of preserving high perceptual quality at extremely low bitrates. While traditional compression standards, e.g., H.266/VVC (VVC), and learning-based neural image compression methods (groupedmixer; jiang2023mlic; 10155660; Cheng_2020_CVPR; balle2018variational; 10510460) primarily optimize for distortion-oriented metrics such as PSNR and MS-SSIM, they often struggle to maintain semantic integrity and visual fidelity under severe bitrate constraints (blau2019rethinking). To address these limitations, recent compression approaches based on generative models have emerged, leveraging powerful generative priors to enable high-fidelity reconstruction even in extreme-low-bitrate scenarios (finetuneVQGAN; illm; hific).
@@
Recent advances in generative models (e.g., tokenizers, GANs, diffusion models) have enabled superior perceptual quality in extreme image compression scenarios. HIFIC (hific) first integrates conditional GANs into VAE frameworks, while MS-ILLM (illm) employs non-binary discriminators for improved statistical fidelity. VQGAN (vqgan) demonstrates strong alignment between latent representations and human perceptual characteristics. Recent works have integrated VQVAE tokenizers into compression frameworks, with VQ-means (finetuneVQGAN) introducing K-means clustering and UIGC (uigc) achieving better rate-perception performance through transformer-based token regeneration. GLC (GLC) further reduces redundancy through latent space compression. Diffusion-based approaches (RDEIC; yang2024lossy; körber2024percosdopenperceptual; jiang2023multi) have also shown promise in generative compression. However, existing methods rely on 2-D architectures that limit compression efficiency due to spatial redundancy. To address this limitation, we propose a 1-D architecture that encodes images into compact 1-D latent representations, enabling more effective extreme image compression.

### v2

@@ Abstract
Recent advances in extreme image compression have revealed that mapping pixel data into highly compact latent representations can significantly improve coding efficiency. However, most existing methods compress images into 2-D latent spaces via convolutional neural networks (CNNs) or Swin Transformers, which tend to retain substantial spatial redundancy, thereby limiting overall compression performance. In this paper, we propose a novel Mixed RWKV-Transformer (MRT) architecture that encodes images into more compact 1-D latent representations by synergistically integrating the complementary strengths of linear-attention-based RWKV and self-attention-based Transformer models. Specifically, MRT partitions each image into fixed-size windows, utilizing RWKV modules to capture global dependencies across windows and Transformer blocks to model local redundancies within each window. The hierarchical attention mechanism enables more efficient and compact representation learning in the 1-D domain. To further enhance compression efficiency, we introduce a dedicated RWKV Compression Model (RCM) tailored to the structure characteristics of the intermediate 1-D latent features in MRT. Extensive experiments on standard image compression benchmarks validate the effectiveness of our approach. The proposed MRT framework consistently achieves superior reconstruction quality at bitrates below 0.02 bits per pixel (bpp). Quantitative results based on the DISTS metric show that MRT significantly outperforms the state-of-the-art 2-D architecture GLC, achieving bitrate savings of $43.75\%$, $30.59\%$ on the Kodak and CLIC2020 test datasets, respectively. All source code are available at https://github.com/luke1453lh/MRT.
@@ Introduction
The explosive growth of visual data has intensified the demand for advanced image compression techniques capable of preserving high perceptual quality at extremely low bitrates. While traditional compression standards, e.g., H.266/VVC (VVC), and learning-based neural image compression methods (groupedmixer; jiang2023mlic; Cheng_2020_CVPR; balle2018variational; 10155660; 10510460) primarily optimize for distortion-oriented metrics such as PSNR and MS-SSIM, they often struggle to maintain semantic integrity and visual fidelity under severe bitrate constraints (blau2019rethinking). To address these limitations, recent compression approaches based on generative models have emerged, leveraging powerful generative priors to enable high-fidelity reconstruction even in extreme-low-bitrate scenarios (finetuneVQGAN; illm; hific).
@@
With the rapid development of deep learning (10003241; 10608071), generative models have achieved remarkable success in various computer vision tasks. Traditional hybrid coding frameworks  (VVC) and their deep-learning-enhanced variants (li2025ustc; 10856241; li2024object; li2024loop) have significantly improved signal fidelity, yet they still suffer from severe quality degradation at extremely low bitrates, leading to unsatisfactory perceptual experience. Recent advances in generative models (e.g., tokenizers, GANs, diffusion models) have enabled superior perceptual quality in extreme image compression scenarios. HIFIC (hific) first integrates conditional GANs into VAE frameworks, while MS-ILLM (illm) employs non-binary discriminators for improved statistical fidelity. VQGAN (vqgan) demonstrates strong alignment between latent representations and human perceptual characteristics. Recent works have integrated VQVAE tokenizers into compression frameworks, with VQ-means (finetuneVQGAN) introducing K-means clustering and UIGC (uigc) achieving better rate-perception performance through transformer-based token regeneration. GLC (GLC) further reduces redundancy through latent space compression. Diffusion-based approaches (RDEIC; yang2024lossy; körber2024percosdopenperceptual; jiang2023multi) have also shown promise in generative compression. However, existing methods rely on 2-D architectures that limit compression efficiency due to spatial redundancy. To address this limitation, we propose a 1-D architecture that encodes images into compact 1-D latent representations, enabling more effective extreme image compression.
@@ Acknowledgment
Appendix A Appendix1: Model Architecture
Our proposed model comprises two main components: a Mixed RWKV-Transformer (MRT), and a dedicated RWKV compression model (RCM). In this section, we provide a detailed description of the proposed MRT and RCM.
Mixed RWKV-Transformer
We present the detailed architecture specifications of the 1D transform with comprehensive hyperparameters. As depicted in Figure 8, the 1D-transform architecture comprises 1 Bi-RWKV block and 4 ViT blocks, all sharing identical dimensions with $d=1024$ and the same ratio of $r=4$.
RWKV Compression Model
As illustrated in Figure 9, the RWKV compression model utilizes multiple Bi-RWKV blocks to effectively capture both spatial and channel dependencies, with all Bi-RWKV blocks sharing identical dimensions of $d=320$ and a ratio of $r=4$. For the 1D SCCTX component, we adopt the entropy model from TCM (tcm) to model the distribution of the 1D latent features, simply replacing Conv blocks with Bi-RWKV blocks.
Appendix B Appendix2: Experiments
Training Details
We train our model using the Open Image v4 training dataset (openimages) with a carefully designed two-stage training strategy. In the first stage, we employ the AdamW optimizer (adamw) with a weight decay of $1\times 10^{-4}$ and a learning rate of $1\times 10^{-4}$, incorporating 10,000 warmup steps followed by a cosine decay schedule. In the second stage, we maintain identical optimizer settings including learning rate, warmup steps, and decay schedule. To enhance training stability, we introduce the discriminator component only after 20,000 training steps. For both training stages, we utilize a batch size of 4, an exponential moving average (EMA) decay of 0.999, and employ mixed-precision training.
Evaluation Details
Evaluation Metrics
We evaluate the performance of our model on the three widely used datasets: CLIC2020 test set (CLIC2020), Kodak (kodak), and DIV2K vaild set (DIV2K) using the following metrics: LPIPS (LPIPS), DISTS (DISTS), and FSIM (FID). Following HIFIC (hific), we evaluate fid by splitting the images into 256 overlapped patches, noting that fid is not computed on the Kodak dataset as kodak only contains limited number of images. All metrics are computed on the full resolution images of three datasets.
Evaluation Methods
We evaluate the performance of LALIC (lalic), MS-ILLM (illm), and HIFIC (hific) by utilizing the official released checkpoints and fine-tuning their pretrained models to extremely low bitrates. As for RDEIC (RDEIC) and DiffEIC (Diffeic), we employ the official released checkpoints for evaluation. For GLC (GLC), we obtain the evaluation results through personal communication with the authors.
Redundancy Analysis
We present a redundancy analysis comparing 2D representations from MS-ILLM (illm) and 1D representations from MRT in Table 4. Spatial redundancy is measured via cosine similarity and $L_{2}$ distance, while channel redundancy is measured via mean feature correlation (MFC). Results show MRT achieves lower spatial and channel redundancy than MS-ILLM, indicating more compact and information-efficient representations.
Complexity Analysis
We provide the computational complexity analysis of different methods in Table 5. MRT outperforms the 2D architecture method MS-ILLM (illm) on BD-DISTS metrics, despite having more model parameters and higher computational complexity. Additionally, MRT achieves better BD-DISTS performance than DiffEIC (Diffeic) while requiring fewer model parameters and maintaining competitive encoding time with significantly reduced decoding time.
Quantitative Results
To give a more comprehensive evaluation, we provide the quantitative results of MRT on CLIC2020 $768\times 768$ test set (CLIC2020), as shown in Figure 11.
Furthermore, we provide PSNR and MS-SSIM results on Kodak (kodak), CLIC2020, and DIV2K (DIV2K) datasets. As illustrated in Figure 10, MRT outperforms the diffusion-based method DiffEIC on all three datasets in both PSNR and MS-SSIM metrics. Notably, pixel-level distortion is not well aligned with human perception, especially at extremely low bitrates (GLC). Although MRT does not achieve the best pixel-level rate-distortion performance, it still exhibits superior visual quality as discussed in the main paper.
Qualitative Results
In this section, we provide additional visual examples across the CLIC2020 test (CLIC2020) and MS-COCO 30K (lin2014microsoft) datasets. For these examples, MRT achieves pleasing visual quality.
Kodak MS-ILLM (2D) MRT (1D)
MFC ↓ 0.30 0.20
cosine ↓ 0.79 0.30
$L_{2}$ distance ↑ 2.46 27.69
Table 4: Redundancy analysis of 2D representations generated by MS-ILLM (illm) and 1D representations generated by MRT.
Models Kodak DIV2K Params(M)
Enc.(s) Dec.(S) Mem.(G) BD-DISTS Enc.(s) Dec.(S) Mem.(G) BD-DISTS
MS-ILLM $0.035$ $0.030$ $1.051$ $0.00\%$ $0.401$ $0.361$ $5.088$ $0.00\%$ $181.0$
DiffEIC $0.146$ $3.908$ $6.690$ $34.52\%$ $0.910$ $58.875$ $21.237$ $40.90\%$ $1379.50$
MRT(ours) $0.164$ $0.191$ $4.738$ $77.64\%$ $0.771$ $1.056$ $15.730$ $76.43\%$ $906.11$
Table 5: Computational complexity analysis of different methods. All tests are conducted on a single NVIDIA A100 GPU.
Figure 8: Overview of Mixed RWKV-Transformer and hyperparameters of 1-D Transform.
Figure 9: Hyperparameters of the RWKV compression model.
Figure 10: PSNR and MS-SSIM results of different methods on Kodak, CLIC2020 and DIV2K datasets.
Figure 11: LPIPS, DISTS, and FID results of different methods on CLIC2020 dataset ($\times768$ resolution).
Figure 12: Qualitative examples on CLIC2020 dataset (full resolution).
Figure 13: Qualitative examples on CLIC2020 dataset ($\times768$ resolution).
Figure 14: Qualitative examples on MS-COCO 30K dataset.